Diane invested her savings in two investment funds. The amount she invested in Fund A was $8,000 less than the amount she invested in Fund B. Fund A returned a 5% profit and Fund B returned a 4% profit. How much did she invest in Fund B, if the total profit from the two funds together was $1040?

Answer by stanbon(75887)

Diane invested her savings in two investment funds. The amount she invested in Fund A was $8,000 less than the amount she invested in Fund B. Fund A returned a 5% profit and Fund B returned a 4% profit. How much did she invest in Fund B, if the total profit from the two funds together was $1040?
----
Equations:
A = B - 8000
5A +4B = 104000
---------------------------
Multiply thru 1st equation by 5:
5A -5B = -40000
5A + 4B= 104000
-----
Subtract and solve for "B":
9B = 144000
B = $16000 (amt. invested in B)
=========================================
Cheers,
Stan H.
============